OpenStudy (anonymous):

find the x and y intercept of the line given by the equation : 8x+9y=13 x-int=? y-int=?

OpenStudy (dan815):

put it in to y = form yourself this time

OpenStudy (jim766):

y intercept is when x = 0 x intercept is when y = 0 sub those in

OpenStudy (jim766):

8x + 9y = 13 8(0) + 9y = 13 solve for y intercept

OpenStudy (anonymous):

so y=13/9 and x= 13/8?

OpenStudy (jim766):

yep

OpenStudy (anonymous):

thanks!

OpenStudy (dan815):